# Deviations from the Nash equilibrium and emergence of tacit collusion in a two-player optimal execution game with reinforcement learning

### v1

@@ 2 Market impact game setting
Focusing on the special case of two players, in [schied2017state] it is proved that the selling schedule at the unique Nash equilibrium is:
$q^{*(1)}_{t}=\frac{1}{2}(\Sigma_{t}+\Delta_{t})\hskip 56.9055pt\text{and}\hskip 56.9055ptq^{*(2)}_{t}=\frac{1}{2}(\Sigma_{t}-\Delta_{t})$ (7)
where:
$\begin{split}\Sigma_{t}=V_{t}e^{-\frac{\kappa t}{6\alpha}}\frac{\sinh{\left(\frac{(N-t)\sqrt{\kappa^{2}+12\alpha\lambda\sigma^{2}}}{6\alpha}\right)}}{\sinh{\left(\frac{N\sqrt{\kappa^{2}+12\alpha\lambda\sigma^{2}}}{6\alpha}\right)}}\hskip 28.45274pt\text{and}\hskip 28.45274pt\Delta_{t}=V_{t}e^{\frac{\kappa t}{2\alpha}}\frac{\sinh{\left(\frac{(N-t)\sqrt{\kappa^{2}+4\alpha\lambda\sigma^{2}}}{2\alpha}\right)}}{\sinh{\left(\frac{N\sqrt{\kappa^{2}+4\alpha\lambda\sigma^{2}}}{2\alpha}\right)}}\end{split}$ (8)
It can be noticed that the Nash inventory holdings, and thus the trading rates, now depend also on the permanent impact $\kappa$, contrary to what happens in the single agent case studied in [almgren]. The open loop setting of the game excludes that the agents have knowledge of each others inventory holdings and, as said above, they interact only through the permanently-impacted price ${S}_{t}$. The price level does not enter directly into the optimal inventory formula, but the permanent impact and the volatility of the asset on the other hand do, proportionally to agents’ risk aversion.
@@
In this setting, a Nash equilibrium exists and is unique. However, one could wonder if, apart from the Nash equilibrium, there are other solutions that might be non-Nash and collusive or, generally speaking, either better or sub-optimal, in terms of average costs obtained by either agent, when compared to the Nash solution. Motivated by this, we aim at studying how and under which conditions agents interacting in such an environment can adopt manipulative or collusive behaviours that consistently deviate from the Nash equilibrium. We use RL to train the agents to trade optimally in the presence of the other agent and we study the resulting equilibria. The aim is to ascertain whether while trading, sub-optimal or collusive cost profiles for either of the agents are attainable. In order to do so, we focus on the case where two risk neutral agents want to liquidate the same initial portfolio made by an amount $q_{0}$ of the same asset. We let the agents play multiple episode instances of the optimal execution problem in a multi-agent market. This means that, defining a trading episode to be the complete unwinding of the initial inventory $q_{0}$ over time $[0,T]$ in $N$ time steps by both agents, the overall game is made by iterated trades at each $t=1,\dots,N$ time-steps and by $B$ iterations of trading, each containing a full inventory liquidation episode. Each iteration is described by two vectors $\vec{v}^{\,(1)}_{(i)}$ and $\vec{v}^{\,(2)}_{(i)}$ containing the trading schedule of the two agents in episode $i$. Moreover, each vector is associated with its IS, according to Eq. 2. We now define the set of admissible strategies and the average IS.
@@
In an iterated game with $B$ independent iterations, the set $\mathcal{A}$ of admissible selling strategies in iteration $i$ is composed by the pair of vectors $\vec{v}^{\,(1,2)}_{(i)}=(\vec{v}^{\,(1)}_{(i)},\vec{v}^{\,(2)}_{(i)})$, such that for $k=1,2$:
  • $\sum^{N}_{t=1}v^{(k)}_{(i),t}=q_{0}$
  • $\vec{v}_{(i),t}^{(k)}$ is $({\mathcal{F}}_{t})_{t\geq 0}$ adapted and progressively measurable
  • $\sum^{N}_{t=1}(v^{(k)}_{(i),t})^{2}<\infty$
For any $\vec{v}^{\,(1,2)}_{(i)}\in\mathcal{A}$ the average Implementation Shortfall is defined as:
$\bar{IS}(\vec{v}^{\,(1,2)}_{(i)})=\frac{1}{N}\sum^{N}_{t=1}IS(\vec{v}^{\,(1,2)}_{(i)},t)=IS(\vec{v}^{\,(1)}_{(i)})+IS(\vec{v}^{\,(2)}_{(i)})$ (9)
Leveraging on [cont2022dynamics] and [cont2021interbank], we define collusive strategies and we show that they are necessarily Pareto-optimal.
@@
In an iterated game with $B$ independent iterations, a pair of vectors of selling schedules $\vec{v}^{\,(1,2)}_{c}=(\vec{v}^{\,(1)}_{c},\vec{v}^{\,(2)}_{c})\,\,\in\mathcal{A}$ is defined to be a collusion if for each iteration $i$ and $\forall\,\,\vec{v}^{\,(1,2)}_{(i)}\in\mathcal{A}$:
$\bar{IS}(\vec{v}^{\,(1,2)}_{c})\leq\bar{IS}(\vec{v}^{\,(1,2)}_{(i)})$
@@
In an iterated game with $B$ independent iterations, a pair of vectors of selling schedules $\vec{v}^{\,(1,2)}_{p}=(\vec{v}^{\,(1)}_{p},\vec{v}^{\,(2)}_{p})\,\,\in\mathcal{A}$ is a Pareto optimal strategy if and only if there does not exist $\vec{v}^{\,(1,2)}_{(i)}\in\mathcal{A}$ such that:
$\begin{split}\forall\,\,i=1,\dots,M\,,\,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ &\bar{IS}(\vec{v}^{\,(1,2)}_{p})\geq\bar{IS}(\vec{v}^{\,(1,2)}_{(i)})\\ \exists\,\,j=1,\dots,M\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ &\bar{IS}(\vec{v}^{\,(1,2)}_{p})>\bar{IS}(\vec{v}^{\,(1,2)}_{(j)})\\ \end{split}$ (10)
@@
By contradiction, suppose that $\vec{v}^{\,(1,2)}_{c}$ is not Pareto optimal, then there must exist an iteration $i$ and a strategy $\vec{u}^{\,(1,2)}_{(i)}=(\vec{u}^{\,(1)}_{(i)},\vec{u}^{\,(2)}_{(i)})$ such that:
$\bar{IS}(\vec{u}^{\,(1,2)}_{(i)})\leq\bar{IS}(\vec{v}^{\,(1,2)}_{c})$ (11)
@@
Having shown that a collusive strategy is in fact the Pareto-optimal strategy, we aim at finding the set of all Pareto-efficient strategies, i.e. those selling strategies that result in IS levels where it is impossible to improve trading costs for one agent without deteriorating the one of the other agent. These strategies are defined to be collusions in this game setup. We define the set of Pareto solutions and, within this set and we find the Pareto-optimum as the minimum within the set of solutions.
@@
For a two-player game, with $B$ independent iterations and risk neutral players that aim to solve the optimal execution problem in a market model defined as in Eq. (4), the Pareto-efficient set of strategies is the set of strategies $\vec{v}^{\,(1,2)}$ such that the IS of one agent cannot be improved without increasing the IS value of the other agent.
We now provide the conditions that allows to find the set of Pareto-efficient solutions.
@@
$\begin{cases}\min_{\vec{v}^{\,(1,2)}}\vec{f}\left(\vec{v}^{\,(1,2)}\right)\\ \text{s.t.}\,\,\{\vec{v}^{\,(1,2)}_{(i)}\in\mathcal{A},\,\,\vec{g}(\vec{v}^{\,(1,2)})=\vec{0}\}\end{cases}$ (P1)
@@
Problem (P1) is a convex multi-objective optimisation problem with $n=N$ design variables $(\vec{v}^{\,(1,2)})$, $k=2$ objective functions, and $m=2$ constraint functions. Leveraging on [gobbi2015analytical], in order to find the Pareto-efficient set of solutions, Problem (P1) can be restated in terms of Fritz-Johns conditions. We define the $\mathbf{L}\in{\mathbb{R}}^{(n+m)\times(k+m)}$ matrix as:
@@
Leveraging on [gobbi2015analytical], a solution $\vec{v}^{\,(1,2),*}$ to the Problem (P2) is a non-trivial Pareto-efficient solution if:
@@
For a two-player game with $B$ independent iterations and risk-neutral players, the Pareto optimal strategy for all iterations $i$ defined as the solution of the problem:
@@
is $\forall\,\,\,t=1,\dots,N$ and $\forall\,i=1,\dots M$:
@@
Considering just an iteration $i$ of the two players’ game, we define $\vec{v}^{\,(1)}=\operatorname*{arg\,min}_{\vec{v}^{\,(1)}}IS(\vec{v}^{\,(1)}|\vec{v}^{\,(2)})$, meaning that the strategy for agent $1$ is a function of the strategy of agent $2$. And thus $\vec{v}^{\,(1)}$ the strategy that minimises the IS of agent $1$ given the strategy of agent $2$. The same consideration holds for agent $2$, where $\vec{v}^{\,(2)}=\operatorname*{arg\,min}_{\vec{v}^{\,(2)}}IS(\vec{v}^{\,(2)}|\vec{v}^{\,(1)})$.
@@
we then set up a constrained optimisation problem, where the constraint binds the agents to just sell their initial inventory $q_{0}$ over the time window considered:
$\begin{split}&\nabla_{\vec{v}^{\,(1,2)}}\left[{F}\left(\vec{v}^{\,(1,2)}\right)+\lambda_{1}\left(\sum^{N}_{t=1}v^{(1)}_{t}-q_{0}\right)+\lambda_{2}\left(\sum^{N}_{t=1}v^{(2)}_{t}-q_{0}\right)\right]=0\\ \end{split}$ (18)
@@
To study whether Collusive Pareto-optima or more generally non-Nash equilibria arise in this setting, we model the two agents by equally and simultaneously training them with RL based on Double Deep Q Learning (DDQL) algorithm where two agents interact in an iterative manner. We then analyse the results under the light of the collusion strategies derived above. The aim is to understand how and which equilibria will eventually be attained, and whether in attaining an equilibrium they tacitly collude in order to further drive the costs of their trading down. We will analyse their interactions in some limiting cases in order to study if and how the equilibrium found changes and adapts to the market dynamics in a model agnostic setting such as the DDQL algorithm.
@@ 3 Double Deep Q Learning for multi-agent impact trading
We model the algorithmic agents by using RL based on Double Deep Q-Learning (DDQL). The setup is similar to the single agent algorithm as in [ning2021double] and in [macrì2024reinforcement], but now we consider two agents interacting in the same environment. Each agent employs two neural networks, namely the main Q-net ($Q_{\text{main}}$) for action selection and the target Q-net ($Q_{\text{tgt}}$) for state evaluation. The four nets are exactly the same and updated at exactly the same rate. The only difference between the two agents is the timing at which they act and the time at which their nets are updated. In fact, in a given time-step, the agent trading as second pays the price impact generated by the first one. To make the game symmetric, similarly to [schied2019market], at each time step a coin toss decides which agent trades first. This ensures the symmetry and guarantees that, as the game unfolds, no advantage in terms of trading timing is present for either of the two agents.
We divide the overall numerical experiment into a training and a testing phase. In the training phase we train both the agents to solve the optimal execution problem. In the testing phase, we employ what we learned in terms of $Q_{\text{main}}$ weights, letting the trained agents play an iterated trading game. All the results shown in Section 4 are obtained in the testing phase.
@@ 3.1 Setting of the numerical experiments
We consider two risk neutral agents whose goal is to unwind an initial position of $q_{0}=100$ shares with initial value $S_{0}=10\$$ within a time window $[0,T]$. The window is divided into $N=10$ time-steps of length $T/N=\tau$. The mid-price $S_{t}$ evolves as in Eq. (4), and the agents sell their whole inventory during the considered time-window. This is called an iteration and in order to train the DDQL algorithm we consider a number of $C$ iterations. This is called a run. Thus, a run of the game is defined to be an iterated game over both $N$ time-steps and in $C$ trading iterations.
Over the $C$ iterations, each agent learns how to trade via an exploration-exploitation scheme, thus using $\epsilon$-greedy policies. This is obtained by changing the weights of their $Q_{\text{main}}$ net in order to individually choose the best policies in terms of rewards, related to the obtained IS. The scheme of the algorithm is thus symmetric, and for each agent it is divided into two phases: an exploration and an exploration phase managed by the parameter $\epsilon\in(0,1]$, that is common to both the agents, that will decrease geometrically during training and globally initialised as $\epsilon=1$. Depending on the phase, the way in which the quantities to sell $v_{t}$ are chosen changes, and it does so for each agent symmetrically. When the agents are exploring, they will randomly select quantities to sell $v_{t}$ in order to explore different states and rewards in the environment, alternatively, when the agents are exploiting, they will use their $Q_{\text{main}}$ net to select $v_{t}$.

### v2

@@ 2 Market impact game setting
Focusing on the special case of two players, in [undefam] it is proved that the selling schedule at the unique Nash equilibrium is:
$q^{(1)^{*}}_{t}=\frac{1}{2}(\Sigma(t)+\Delta(t))\hskip 56.9055pt\text{and}\hskip 56.9055ptq^{(2)^{*}}_{t}=\frac{1}{2}(\Sigma(t)-\Delta(t))\,$, (7)
with:
$\begin{split}\Sigma(t)=Qe^{-\frac{\kappa t}{6\alpha}}\frac{\sinh{\left(\frac{(N-t)\sqrt{\kappa^{2}+12\alpha\lambda\sigma^{2}}}{6\alpha}\right)}}{\sinh{\left(\frac{N\sqrt{\kappa^{2}+12\alpha\lambda\sigma^{2}}}{6\alpha}\right)}}\hskip 28.45274pt\text{and}\hskip 28.45274pt\Delta(t)=\tilde{Q}e^{\frac{\kappa t}{2\alpha}}\frac{\sinh{\left(\frac{(N-t)\sqrt{\kappa^{2}+4\alpha\lambda\sigma^{2}}}{2\alpha}\right)}}{\sinh{\left(\frac{N\sqrt{\kappa^{2}+4\alpha\lambda\sigma^{2}}}{2\alpha}\right)}}\end{split}$ (8)
where $Q=q^{(1)}_{0}+q^{(2)}_{0}$ and $\tilde{Q}=q^{(1)}_{0}-q^{(2)}_{0}$. It can be noticed that the Nash inventory holdings, and thus the trading rates, now depend also on the permanent impact $\kappa$, contrary to what happens in the single agent case studied in [undefb]. The price level does not enter directly into the optimal inventory formula, but the permanent impact and the volatility of the asset on the other hand do, proportionally to agents’ risk aversion.
Remark 1.
In [undefam] the authors derive the unique Nash equilibrium considering the continuous time version of Eq.(4). We will use the Nash inventory holdings and the selling strategy associated to Eq.(7)-(8) discretised over time. We choose to consider the discretised version of the Nash equilibrium, and to derive results in discrete time as they are functional to the numerical and deep RL experiments, as well as to highlight the differences that might arise when autonomous agents are trained together in a similar but subtly different environment. The Nash equilibrium in the discrete time setting was derived in Sec. 2 of [undefu], considering a constant decay kernel in a transient impact market. [^8] In [undefak]the authors show that the high-frequency limits of the discrete-time equilibrium costs converge to the expected costs in the continuous-time Nash equilibrium.
@@
In this setting, as shown a unique Nash equilibrium exists but, leveraging on reviewed literature, other non-Nash equilibria might as well. Especially, supra-competitive and collusive equilibria have been shown to exists in oligopolistic settings such as the one considered in a market impact game. To this end, we aim at studying the behaviour of two agents who have to solve an optimal execution problem in a multi-agent setting. We study the problem from a learning perspective, meaning that we will use deep RL to train the agents to solve the liquidation problem in such setup; we then study the strategies found by the agents, comparing the cost of the learnt strategies with the ones prescribed by both the Nash equilibrium strategy and the Pareto optimal strategy, that will be shown to be the collusive strategy. In order to do so, we focus on the case where two risk neutral agents want to liquidate the same initial portfolio made by an amount $q_{0}$ of the same asset. In the numerical setup we let the agents play multiple episode instances of the optimal execution problem. This means that, defining a trading episode to be the complete unwinding of the initial inventory $q_{0}$ over time $[0,T]$ in $N$ time steps by both agents, the overall game is made by repeated trades at each $t=1,\dots,N$ time-steps and by $B$ iterations of trading, each containing a full inventory liquidation episode. Therefore, for each iteration $i$ we will consider two vectors $\vec{v}^{\,(1)}$ and $\vec{v}^{\,(2)}$ containing the trading schedule for the two agents, .
The numerical setup is justified by the episodical nature of the game under consideration, as by definition -and differently from other setups- the optimal execution problem happens over a finite time-window. Each vector is associated with its IS cost, according to Eq. 2. This cost, along with the strategies’ vectors are going to be the main objective of our study. We now define the set of admissible strategies and the average IS cost.
@@
The set $\mathcal{A}$ of admissible selling strategies is composed by the pair of vectors $\vec{v}^{\,(1,2)}=(\vec{v}^{\,(1)},\vec{v}^{\,(2)})$, such that for $k=1,2$:
  • $\sum^{N}_{t=1}v^{(k)}_{{t}}=q_{0}$
  • $\vec{v}_{{t}}^{(k)}$ is $({\mathcal{F}}_{t})_{t\geq 0}$ - adapted and progressively measurable
  • $\sum^{N}_{t=1}(v^{(k)}_{{t}})^{2}<\infty$
For any $\vec{v}^{\,(1,2)}\in\mathcal{A}$ the average Implementation Shortfall is defined as:
$\bar{IS}\left(\vec{v}^{\,(1,2)}\right)=\frac{1}{N}\sum^{N}_{t=1}IS\left(\vec{v}^{\,(1,2)}_{{t}}\right)\,$. (9)
where $IS\left(\vec{v}^{\,(1,2)}_{{t}}\right)=IS\left(\vec{v}^{\,(1)}_{{t}}\right)+IS\left(\vec{v}^{\,(2)}_{{t}}\right)$.
Leveraging on [undefs] and [undefr], we define collusive strategies and we show that they are necessarily Pareto-optimal. Notice that, given the iterated nature of the game due to the time and inventory constraints that -by design- constrain the agents, we will consider the following equilibrium properties to hold over several iterations of the game. Although the same properties would hold in a one-shot game, we aim at studying what equilibrium would naturally arise when the agents are simultaneously learning, without being told to either cooperate or compete. The theoretical results stated below are functional to the analysis of the learned equilibria, therefore in the numerical experiments part we depart from a one-shot game setup in favour of an iterated game setup.
@@
A pair of vectors of selling schedules $\vec{v}^{\,(1,2)}_{c}=(\vec{v}^{\,(1)}_{c},\vec{v}^{\,(2)}_{c})\,\,\in\mathcal{A}$ is defined to be a collusion if $\forall\,\,\vec{v}^{\,(1,2)}\in\mathcal{A}$:
$\bar{IS}(\vec{v}^{\,(1,2)}_{c})\leq\bar{IS}(\vec{v}^{\,(1,2)})$
@@
A pair of vectors of selling schedules $\vec{v}^{\,(1,2)}_{p}=(\vec{v}^{\,(1)}_{p},\vec{v}^{\,(2)}_{p})\,\,\in\mathcal{A}$ is a Pareto optimal strategy if and only if there does not exist $\vec{v}^{\,(1,2)}\in\mathcal{A}$ such that:
[…]
@@
By contradiction, suppose that $\vec{v}^{\,(1,2)}_{c}$ is not Pareto optimal, then there must exist an iteration $i$ and a strategy $\vec{u}^{\,(1,2)}=(\vec{u}^{\,(1)},\vec{u}^{\,(2)})$ such that:
$\bar{IS}(\vec{u}^{\,(1,2)})\leq\bar{IS}(\vec{v}^{\,(1,2)}_{c})$ (11)
@@
Having shown that a collusive strategy is in fact the Pareto-optimal strategy, we aim at finding the set of all Pareto-efficient strategies, i.e. those selling strategies that result in IS levels where it is impossible to improve trading costs for one agent without deteriorating the one of the other agent. In other words, we are looking for those strategies where an agent is better off only if the other agent is be worse off. We define the set of Pareto solutions and, within this set and we find the Pareto-optimum as the minimum within the set of solutions.
@@
For the two-player game, considering risk neutral players that aim to solve the optimal execution problem in a market model defined as in Eq. (4), the Pareto-efficient set of strategies is the set of strategies $\vec{v}^{\,(1,2)}$ such that the IS of one agent cannot be improved without increasing the IS value of the other agent.
We now provide the conditions that allow to find the set of Pareto-efficient solutions as a minimisation of the Implementation Shortfall functional of both the agents. The minimisation has to be intended in the sense of Pareto dominance component-wise.
@@
$\begin{cases}\min_{\vec{v}^{\,(1,2)}}\vec{f}\left(\vec{v}^{\,(1,2)}\right)\\ \text{s.t.}\,\,\{\vec{v}^{\,(1,2)}\in\mathcal{A},\,\,\vec{g}(\vec{v}^{\,(1,2)})=\vec{0}\}\end{cases}$ (P1)
@@
Problem (P1) is a convex multi-objective optimisation problem with $n=N$ design variables $(\vec{v}^{\,(1,2)})$, $k=2$ objective functions, and $m=2$ constraint functions. Leveraging on [undefx], in order to find the Pareto-efficient set of solutions, Problem (P1) can be restated in terms of Fritz-Johns conditions. We define the $\mathbf{L}\in{\mathbb{R}}^{(n+m)\times(k+m)}$ matrix as:
@@
Leveraging on [undefx], a solution $\vec{v}^{\,(1,2),*}$ to the Problem (P2) is a non-trivial Pareto-efficient solution if:
@@
For a two-player game with risk-neutral players, the Pareto optimal strategy is the solution of the problem:
@@
is $\forall\,\,\,t=1,\dots,N$:
@@
First define $\vec{v}^{\,(1)}=\operatorname*{arg\,min}_{\vec{v}^{\,(1)}}IS(\vec{v}^{\,(1)}|\vec{v}^{\,(2)})$, meaning that the strategy for agent $1$ is a function of the strategy of agent $2$. And thus $\vec{v}^{\,(1)}$ the strategy that minimises the IS of agent $1$ given the strategy of agent $2$. The same consideration holds for agent $2$, where $\vec{v}^{\,(2)}=\operatorname*{arg\,min}_{\vec{v}^{\,(2)}}IS(\vec{v}^{\,(2)}|\vec{v}^{\,(1)})$.
@@
we then set up a constrained optimisation problem, where the constraint binds the agents to just sell their initial inventory $q_{0}$ over the time window considered , where now $\lambda_{1},\lambda_{2}$ are multipliers:
[…]
@@
To study whether collusive Pareto-optima, or more generally supra-competitive non-Nash equilibria, arise in this setting, we model the two agents by equally and simultaneously training them with Double Deep Q Learning (DDQL), a deep RL algorithm. We then analyse the results under the light of the collusion strategies derived above. The aim is to understand how and which equilibria will eventually be learned, and whether in attaining an equilibrium they tacitly collude in order to further drive the costs of their trading down, sustained by a punishment mechanism.
We choose to implement this algorithm and use it for our analysis for two main reasons. On the one hand, using Double Deep Q-Learning allows us to be more consistent with the extant literature that mainly uses Q-learning, outlining if observing tacit-collusive behaviours is still possible when using deep neural networks and memory of past states. In fact, from the literature it is well known that such an off-policy algorithm works well in the single agent case (see, for instance, [undefaf, undefai]), making it a safe and sensible ground choice when it comes to extend the framework to more than one agent. On the other hand, and more methodologically, we have chosen Double Deep Q-learning because this algorithm directly maximizes the expected discounted rewards without explicitly parametrizing the policy. Thus, this algorithm choice allows us to analyse the problem from a purely reward driven point of view, as the Q-values of the agents are the quantities that are optimised through the use of neural networks without the need to visit every state. We think that an interesting extension of our work could be to compare different RL algorithms to identify pros and cons of each of them.
@@ 3 Double Deep Q Learning for multi-agent impact trading
We model the algorithmic agents by using RL based on Double Deep Q-Learning (DDQL). The setup is similar to the single agent algorithm as in [undefai] and in [undefaf], but now we consider two agents interacting in the same environment. The agents are modelled to sequentially interact at each time step to trade either as first or as second. In principle, the game does not need to be a sequential move game but as in real markets, agents cannot trade simultaneously and for this reason we have chosen this numerical setup for the experiments. Moreover, the choice of the algorithm is motivated by the fact that this algorithm has been proven to work with both real data, as in [undefai], and especially with simulated data where the permanent impact and the risk aversion parameter are present as in [undefaf].
Each agent employs two neural networks, namely the main Q-net ($Q_{\text{main}}$) for action selection and the target Q-net ($Q_{\text{tgt}}$) for state evaluation. The four nets share the same architecture, their weights evolve independently during the training phase and they are updated at exactly the same rate. The only difference between the two agents is the timing at which they act and the time at which their nets are updated. In fact, in a given time-step, the agent trading as second pays the price impact generated by the first one. To make the game symmetric, as specified above and similarly to [undefal], at each time step a coin toss decides which agent trades first. This ensures the symmetry and guarantees that, as the game unfolds, no advantage in terms of trading timing is present for either of the two agents.
We divide the overall numerical experiment into a training and a testing phase. In the training phase we train both the agents to solve the optimal execution problem. In the testing phase, we employ what learnt in terms of $Q_{\text{main}}$ weights, letting the trained agents play an iterated trading game. All the results shown in Section 4 are obtained in the testing phase.
@@ 3.1 Setting of the numerical experiments
We consider two risk neutral agents whose goal is to unwind an initial position of $q_{0}=100$ shares with initial value $S_{0}=10\mathdollar$ within a time window $[0,T]$. The window is divided into $N=10$ time-steps of length $T/N=\tau$. The mid-price $S_{t}$ evolves as in Eq. (4), and the agents sell their whole inventory during the considered time-window. This is called an iteration and in order to train the DDQL algorithm we consider a number of $C$ iterations. This is called a run. Thus, a run of the game is defined to be a repeated game over both $N$ time-steps and in $C$ trading iterations.
Over the $C$ iterations, each agent learns how to trade via an exploration-exploitation scheme, thus using $\epsilon$-greedy policies. This is obtained by changing the weights of their $Q_{\text{main}}$ net in order to individually choose the best policies in terms of rewards, related to the obtained IS. The scheme of the algorithm is thus symmetric, and for each agent it is divided into two ‘phases’: an exploration and an exploitation phase managed by the parameter $\epsilon\in(0,1]$, that is common to both the agents [^9] More practically, the agents will explore with probability $\epsilon$ and exploit with probability $1-\epsilon$., that will decrease geometrically during training and globally initialised as $\epsilon=1$. Depending on the phase, the way in which the quantities to sell $v_{t}$ are chosen changes, and it does so for each agent symmetrically. When the agents are exploring, they will randomly select quantities to sell $v_{t}$ in order to explore different states and rewards in the environment, alternatively, when the agents are exploiting, they will use their $Q_{\text{main}}$ net to select $v_{t}$.